THE AMMONIA DIMER REVISITED

RICHARD DAWES, Missouri University of Science and Technology, Rolla, MO 65409-0010; \underline{AD VAN DER AVOIRD}, Radboud University, 6525 AJ Nijmegen, The Netherlands.

The conclusion from microwave spectra by Nelson, Fraser, and Klemperer ^{a} that the ammonia dimer has a nearly cyclic structure led to much debate about the issue of whether (NH_{3})_{2} is hydrogen bonded. This structure was surprising because most ab initio calculations led to a classical, nearly linear, hydrogen-bonded structure. An obvious explanation of the discrepancy between the outcome of these calculations and the microwave data which led Nelson et al. to their “surprising structure” might be the effect of vibrational averaging: the electronic structure calculations focus on finding the minimum of the intermolecular potential, the experiment gives a vibrationally averaged structure. Isotope substitution studies seemed to indicate, however, that the complex is nearly rigid. Additional data became available from high-resolution molecular beam far-infrared spectroscopy in the Saykally group ^{b}. These spectra, displaying large tunneling splittings, indicate that the complex is very floppy. The seemingly contradictory experimental data were explained when it became possible ^{c} to calculate the vibration-rotation-tunneling (VRT) states of the complex on a six-dimensional intermolecular potential surface. The potential used was a simple model potential, with parameters fitted to the far-infrared data. Now, for the first time, a six-dimensional potential was computed by high level ab initio methods and this potential will be used in calculations of the VRT states of (NH_{3})_{2} and (ND_{3})_{2}. So, we will finally be able to answer the question whether the conclusions from the model calculations are indeed a valid explanation of the experimental data.

aD. Nelson, G. T. Fraser, and W. Klemperer J. Chem. Phys. \underline{83} 6201 (1985)

bJ. G. Loeser, C. A. Schmuttenmaer, R. C. Cohen, M. J. Elrod, D. W. Steyert, R. J. Saykally, R. E. Bumgarner, and G. A. Blake J. Chem. Phys. \underline{97} 4727 (1992)

cE. H. T. Olthof, A. van der Avoird, and P. E. S. Wormer J. Chem. Phys. \underline{101} 8430 (1994); E. H. T. Olthof, A. van der Avoird, P. E. S. Wormer, J. G. Loeser, and R. J. Saykally J. Chem. Phys. \underline{101} 8443 (1994)